# LEREL: Lipschitz Continuity-Constrained Emotion Recognition Ensemble Learning For Electroencephalography

### v1

@@ Abstract
Accurate and efficient perception of emotional states in oneself and others is crucial, as emotion-related disorders are associated with severe psychosocial impairments. While electroencephalography (EEG) offers a powerful tool for emotion detection, current EEG-based emotion recognition (EER) methods face key limitations: insufficient model stability, limited accuracy in processing high-dimensional nonlinear EEG signals, and poor robustness against intra-subject variability and signal noise. To address these challenges, we propose LEREL (Lipschitz continuity-constrained Emotion Recognition Ensemble Learning), a novel framework that significantly enhances both the accuracy and robustness of emotion recognition performance. The LEREL framework employs Lipschitz continuity constraints to enhance model stability and generalization in EEG emotion recognition, reducing signal variability and noise susceptibility while maintaining strong performance on small-sample datasets. The ensemble learning strategy reduces single-model bias and variance through multi-classifier decision fusion, further optimizing overall performance. Experimental results on three public benchmark datasets (EAV, FACED and SEED) demonstrate LEREL’s effectiveness, achieving average recognition accuracies of 76.43%, 83.00% and 89.22%, respectively.
Keywords Electroencephalography (EEG)  $\cdot$ Lipschitz Continuity  $\cdot$ Ensemble Learning  $\cdot$ Emotion Recognition
@@ 1 Introduction
Emotion recognition – the capacity to assess emotional states in oneself and others – is fundamental to social functioning. This critical ability is impaired in conditions such as autism spectrum disorder, depression and schizophrenia, where deficits are strongly associated with adverse psychosocial outcomes including social withdrawal, self-harm behaviors, and suicidal ideation. The perceptual efficiency and accuracy of emotion recognition therefore represent vital targets for both clinical intervention and computational modeling approaches. Electroencephalography (EEG)-based emotion recognition (EER) technology shows great potential for mental health applications, enabling objective assessment and diagnosis of emotional disorders to facilitate early intervention and precision treatment strategies [1] [2]. Emotions are complex, dynamic psychological states shaped by interpersonal and neurobiological factors. The study of emotions poses significant research challenges, particularly in emotion recognition. While no universal consensus exists on emotion categorization, individual differences in emotional expression, especially among individuals with autism and depression, are critically important for both theoretical and clinical applications [3]. This underscores the need for objective emotion assessment tools capable of precisely capturing internal emotional states to inform clinical diagnosis and guide personalized therapeutic interventions.
Recent advances in EER have focused on developing more accurate [4] [5] and robust [6] models. Promising approaches include multimodal data fusion of EEG with other physiological signals [7] [8] [9] [10] [11] [12] [13] and the construction of brain network features to capture complex signal patterns [14]. Despite these innovations, EER still faces significant challenges, including poor signal-to-noise ration, susceptibility to artifacts, high intra-individual variability [15] [16] [17] [18] [19], and limitations in model generalizability.
Lipschitz continuous constraints provide a rigorous mathematical framework to enhance model stability by restricting gradient variations, ensuring outputs remain robust to minor perturbations. For example, Jia et al. demonstrated improved node-level adversarial robustness in graph neural networks through Lipschitz regularization, significantly enhancing stability with noisy data [20]. Their subsequent work applied similar constraints in a fairness regularization method to control output variations and ensure individual fairness [21]. These constraints also strengthen generalization by regularizing model behavior across diverse data distributions and individual variations. This makes Lipschitz continuous constraints particularly valuable for addressing EER challenges, where signal variability and noise sensitivity are persistent issues [22] [23] [21].
Figure 1: Overall framework of the LEREL
@@
In this study, to address these challenges in EER, we propose a Lipschitz continuity-constrained Emotion Recognition Ensemble Learning (LEREL) framework with the following key contributions:
@@
  • The integration of Lipschitz continuity constraints enhances model stability and generalization during training, effectively addressing EEG signal variability while improving robustness and recognition accuracy.
@@
  • The LEREL framework employs an ensemble learning strategy that harnesses the synergistic strengths of multiple classifiers, significantly enhancing both the precision and resilience of emotion recognition.
@@
  • The LEREL framework demonstrates outstanding performance in emotion recognition across multiple benchmark datasets, confirming its effectiveness and generalizability.
@@ 2.1 Problem Definition
Generally speaking, the EEG signal can be expressed as
$\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$ (1)
where $\mathbf{x}_{T}\in\mathbb{R}^{C}$ represents the signal vector at time point $t$, $C$ is the number of EEG channels, and $T$ is the length of the time series. Further, we can define the emotional category set as following:
$\mathcal{Y}=\{y_{1},y_{2},\dots,y_{K}\}$ (2)
where $K$ is the total number of emotional categories. Thus, we can define the EER problem as a task that aims to discern an individual’s emotional state at a particular moment or over a temporal span based on EEG data. The corresponding mathematic expression can formulated as the emotional category label $\hat{y}$ prediction process based on the given $\mathbf{X}$, following the maximization of the prediction probability as:
@@ 2.3 Lipschitz Gradient-Constrained Band Extraction Component (LGCBE)
The Lipschitz Gradient-Constrained Band Extraction (LGCBE) component is an advanced neural network module designed to process complex signal data. It integrates band decomposition, attention mechanisms, and Lipschitz constraints to enhance signal quality and extract key features. By capturing higher-order relationships and dynamically adjusting modality weights, it effectively handles complex interactions and contextual nuances, improving the accuracy and robustness of signal processing tasks.
Figure 2: The implementation of LGCBE
@@
The resulting frequency-domain signal $\mathbf{F}$ has dimensions with $[B,C,F]$, where $F$ is the dimension of the frequency domain. For each frequency band $b$ (defined according to standard EEG frequency bands such as delta, theta, alpha, beta, and gamma), the band mask is computed and band features are extracted as:
@@
where $\text{MLP}_{\text{spectral}}$ is a network comprising two fully connected layers.
To satisfy the Lipschitz continuity constraints, the spectral attention scores are normalized:
@@
where $\mathbf{W}_{s,b}$ is the expanded form of the spectral weights. The signal is transformed back to the time domain using the Inverse Fast Fourier Transform (IFFT):
$\mathbf{y}=\text{IFFT}(\mathbf{X}_{\text{combined}},n=T,\text{dim}=-1)$ (14)
and normalization is applied:
$\mathbf{y}=\text{LayerNorm}(\mathbf{y})$ (15)
To enhance model stability, the spectral normalization is applied to the weights of the linear layers:
@@ 2.4 Lipschitz Gradient-Constrained Attention Component (LGCA)
In recent years, the attention mechanism has made significant progress in the field of deep learning, especially in handling sequential data and graph-structured data. The Transformer architecture has achieved efficient parallel computation and long-range dependency modeling through self-attention mechanisms [28]. However, the standard dot-product self-attention mechanism tends to exhibit overconfidence when dealing with unbounded input domains, meaning that the model provides high-confidence predictions even when it is uncertain [29]. Moreover, deep Graph Attention Networks (GAT) are prone to gradient explosion when increasing the number of layers, which limits the model’s performance in tasks involving long-range dependencies [30].
To address these issues, researchers have proposed several improvements. For example, Ye et al. introduced Lipschitz regularization to mitigate the overconfidence of Transformer models [29]. Dasoulas et al. proposed a Lipschitz normalization method for self-attention layers, significantly improving the performance of deep Graph Neural Networks [30]. Lipschitz continuity regularization enhances neural networks’ robustness and generalization ability while stabilizing the training process [31, 32]. Inspired by these studies, we designed a Lipschitz Gradient-Constrained Attention component (LGCA) to enhance the model’s stability, robustness, and generalization ability by restricting its Lipschitz constant [33].
@@
where $\mathbf{S}$ denote the scaled dot-product scores and $d_{h}$ is the dimension of each attention head.
The attention scores $\mathbf{S}$ are normalized to a probability distribution using the Softmax function as follow:
$\mathbf{P}=\text{softmax}(\mathbf{S})$ (19)
where $\mathbf{P}$ denote the resulting attention weights. Then, the dropout operation is applied to the attention weights $\mathbf{P}$ to prevent overfitting:
$\mathbf{P}^{\prime}=\text{dropout}(\mathbf{P})$ (20)
Next, the value tensor $\mathbf{V}$ is combined with the attention weights $\mathbf{P}\textquoteright$ ’ through weighted summation, and the multi-head results are merged into the final output through the output projection layer. Let $\mathbf{Y}$ denote the final output:
@@
Figure 3: The implementation of LGCA
Figure 4: The implementation of LGCN
@@ 2.5 Lipschitz Gradient-Constrained Normalization Component (LGCN)
This step ensures that the gradient changes of $\mathbf{X}_{\text{lip}}$ are bounded by the Lipschitz constant $L$, which improves training stability and enhances the model’s generalization capability when processing complex EEG signals.
@@ 3.1 Datasets
To comprehensively evaluate our LEREL, we utilized three representative EEG-based emotion recognition datasets.
The EAV dataset [34] provides a multimodal resource for affective computing research, combining synchronized EEG, audio, and video recordings from 42 participants engaged in naturalistic conversational interactions. This dataset captures five distinct emotional states—neutral, anger, happiness, sadness, and calmness—and comprises 8,400 annotated records, establishing itself as a pioneering benchmark in emotion recognition studies.
The FACED dataset [35] advances EEG-based emotion recognition research by leveraging 32-channel electrophysiological recordings from 123 participants exposed to 28 emotionally evocative video stimuli. This comprehensive resource spans nine discrete emotion categories and is further enriched with subjective self-assessment labels for emotional states, alongside multidimensional affective parameters such as arousal and valence.

### v2

# LEL: A Novel Lipschitz Continuity-constrained Ensemble Learning Model for EEG-based Emotion Recognition

@@ Abstract
The accurate and efficient recognition of emotional states in oneself and others is critical, as impairments in this ability can lead to significant psychosocial difficulties. While electroencephalography (EEG) offers a powerful tool for emotion detection, current EEG-based emotion recognition (EER) methods face key limitations: insufficient model stability, limited accuracy in processing high-dimensional nonlinear EEG signals, and poor robustness against intra-subject variability and signal noise. To address these challenges, we introduce LEL (Lipschitz continuity-constrained Ensemble Learning), a novel framework that enhances EEG-based emotion recognition. By integrating Lipschitz continuity constraints, LEL ensures greater model stability and improves generalization, thereby reducing sensitivity to signal variability and noise while significantly boosting the model’s overall accuracy and robustness. Its ensemble learning strategy optimizes overall performance by fusing decisions from multiple classifiers to reduce single-model bias and variance. Experimental results on three public benchmark datasets (EAV, FACED and SEED) demonstrated the LEL’s state-of-the-art performance, achieving average recognition accuracies of 76.43%, 83.00% and 87.22%, respectively. The official implementation codes are released at https://github.com/NZWANG/LEL.
Keywords Electroencephalography (EEG)  $\cdot$ Lipschitz Continuity  $\cdot$ Ensemble Learning  $\cdot$ EEG-based Emotion Recognition (EER)
@@ 1 Introduction
Emotion recognition—the capacity to assess emotional states—is fundamental to social functioning, enabling relationship building, social understanding, and adaptive responses. While sentiment analysis offers a method for efficient psychological assessment [1, 2, 3], this capacity is often significantly impaired in individuals with neurodevelopmental and psychiatric conditions such as Autism Spectrum Disorder (ASD), Attention-Deficit/Hyperactivity Disorder (ADHD), depression and schizophrenia [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. These impairments lead to substantial personal suffering and societal burdens [14], necessitating efficient and accurate recognition tools for both clinical and computational applications [15].
EEG-based Emotion Recognition (EER) addresses this need by providing an objective assessment of emotional disorders, particularly in cases where communication barriers exist [16]. EEG’s direct measurement of neurophysiological activity captures internal states with high temporal resolution, enabling: 1) detection of rapid emotional dynamics critical for mental health research, 2) non-invasive continuous monitoring for longitudinal studies, and 3) practical deployment through single-channel approaches [17, 18]. This capability enables early intervention and facilitates precision treatment strategies [1].
Emotion complexity—shaped by neurobiological, interpersonal, cultural, and individual factors [2]—presents fundamental challenges. The absence of universal categorization frameworks and significant expression variations in clinical populations [3] necessitate precise quantification of internal states for diagnostics and personalized therapies [19, 20]. Recent EER advancements exemplify sophisticated solutions: FreqDGT [21] and the EEG emotion copilot [22] leverage EEG’s spatiotemporal dynamics to advance objective emotion modeling [23].
Figure 1: (a): Overall framework of the LEL (b): The architecture of LGCBE module (c): The architecture of LGCN module (d): The architecture of LGCA module.
@@
In this study, we propose a Lipschitz continuity-constrained Ensemble Learning (LEL) framework for advancing EER with the following key contributions:
@@
  • The integration of Lipschitz continuity constraints across different modules enhances model stability and generalization during training. This effectively addresses EEG signal variability while improving robustness and recognition accuracy. By restricting the Lipschitz constant, the model’s sensitivity to input perturbations is controlled, thereby enhancing its robustness against noise and artifacts in EEG signals.
@@
  • The LEL employs an ensemble learning strategy that harnesses the synergistic strengths of multiple classifiers, significantly enhancing both the precision and resilience of emotion recognition. By combining the predictions of multiple classifiers, we can reduce the risk of overfitting associated with individual classifiers and improve the model’s ability to recognize complex emotional patterns.
@@
  • The LEL demonstrates outstanding performance in emotion recognition across multiple benchmark datasets, confirming its effectiveness and generalizability.
The remainder of this paper is organized as follows. Section 2 reviews recent advances and key challenges in EER and Lipschitz continuity constraints. Section 3 delineates the methodology of the proposed LEL framework, detailing the integration of Lipschitz constraints and ensemble learning strategies. Section 4 presents the experimental setup and assesses the framework’s performance across multiple benchmark datasets. Section 5 interprets the experimental results and discusses their implications for EER research and applications. Finally, Section 6 summarizes the principal contributions and outlines prospective research directions.
2 Related Work
2.1 Recent Advances and Challenges in EER
Recent advancements in EER have been characterized by the development of increasingly accurate [24, 17] and robust models [25]. A prominent strategy contributing to this progress is multimodal data fusion, which integrates EEG signals with other physiological modalities such as eye movement and heart rate. This approach enables the extraction of complementary information from diverse sources, thereby facilitating a more holistic understanding of emotional states [16]. For instance, Kang et al. introduced a hypergraph multimodal learning framework that effectively combines EEG with additional physiological signals to enhance emotion recognition accuracy [14].
Another significant development in EER is the construction of brain network features to capture complex signal patterns and dynamic neural interactions. This method provides deeper insights into the neural mechanisms underlying emotional processing [26]. The application of two-dimensional convolutional neural networks has shown promise in improving emotion recognition performance [18]. Spatial-temporal transformers further refine these systems by modeling both spatial and temporal dependencies in EEG data [23]. A case-based channel selection technique using interpretable transformer networks optimizes efficiency and accuracy [20]. A novel direction involves neural multimodal contrastive representation learning, applied to EEG-based visual decoding, offering a fresh perspective on emotion recognition methodologies [19].
Despite these promising developments, EER continues to face several critical challenges. Firstly, the inherently low signal-to-noise ratio of EEG data complicates feature extraction, as weak brain signals are often obscured by background noise [26]. Secondly, EEG recordings are highly susceptible to artifacts arising from eye blinks, muscle movements, and other physiological activities, which can distort the data and lead to erroneous interpretations [27]. Thirdly, there exists substantial intra-individual variability in EEG signals, where the brain activity of the same individual may vary significantly across different contexts or time points, posing difficulties for model stability and reliability [28]. Lastly, the generalization capability of current EER models remains limited. Models that perform well on specific datasets often exhibit a marked decline in performance when applied to different datasets or populations, thereby hindering the scalability and practical deployment of EER technologies [29, 30, 31].
2.2 Lipschitz Continuity Constraints and Their Applications
In the field of machine learning, ensuring model stability and robustness is crucial, especially when dealing with noisy or perturbed data. Lipschitz continuous constraints have emerged as a powerful tool to achieve this. They provide a rigorous mathematical framework that enhances model stability by restricting gradient variations. By limiting the rate of change of the output with respect to the input, these constraints ensure the model’s robustness to minor perturbations. This is particularly beneficial in real-world applications where data can be messy and unpredictable. For example, Jia et al. demonstrated improved node-level adversarial robustness in graph neural networks through Lipschitz regularization, significantly enhancing stability with noisy data [32]. Their subsequent work applied similar constraints in a fairness regularization method to control output variations and ensure individual fairness [33].
Lipschitz continuous constraints also strengthen the generalization ability of models. By regularizing model behavior across diverse data distributions and individual variations, these constraints help in creating models that perform consistently across different scenarios. This makes Lipschitz continuous constraints particularly valuable for addressing EER challenges, where signal variability and noise sensitivity are persistent issues [34] [35] [33]. While Lipschitz constraints address model stability, attention mechanisms significantly advance deep learning models’ ability to handle complex data structures. The Transformer architecture achieves efficient parallel computation and long-range dependency modeling through self-attention mechanisms [36]. However, the standard dot-product self-attention mechanism tends toward overconfidence with unbounded input domains, producing high-confidence predictions despite uncertainty [37]. Moreover, deep Graph Attention Networks (GAT) are prone to gradient explosion when increased layers, limiting their performance in long-rang dependency tasks [38].
To address these issues, researchers have proposed several improvements. For example, Ye et al. introduced Lipschitz regularization to mitigate the overconfidence of Transformer models [37]. Dasoulas et al. proposed a Lipschitz normalization method for self-attention layers, significantly improving the performance of deep Graph Neural Networks [38].
@@ 3.1 Problem Definition
Generally speaking, the EEG signal can be expressed as $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$, where $\mathbf{x}_{T}\in\mathbb{R}^{C}$ represents the signal vector at time point $t$, $C$ is the number of EEG channels, and $T$ is the length of the time series. Further, we can define the emotional category set $\mathcal{Y}=\{y_{1},y_{2},\dots,y_{K}\}$, $K$ is the total number of emotional categories. Thus, we can define the EER problem as a task that aims to discern an individual’s emotional state at a particular moment or over a temporal span based on EEG data. The corresponding mathematic expression can be formulated as the emotional category label $\hat{y}$ prediction process based on the given $\mathbf{X}$, following the maximization of the prediction probability as:
@@ 3.3 Lipschitz Gradient-Constrained Band Extraction Component (LGCBE)
By integrating band decomposition, attention mechanisms and Lipschitz constraints, the LGCBE significantly enhances signal quality and precisely extracts key features. The core strength of LGCBE lies in its ability to capture higher-order relationships within the signals while dynamically adjusting modality weights to enable flexible resource allocation based on signal characteristics and context, effectively handling complex interactions and subtle contextual changes. This dynamic adjustment mechanism not only increases the flexibility of signal processing but also enhances the robustness of the system, enabling it to maintain high performance even in the presence of noise interference or incomplete data. The implementation of LGCBE is shown in Figure 1b, while LGCBE-E, LGCBE-F, and LGCBE-B represent the three different output results of the module, corresponding to channel energy, channel feature and band feature, respectively. These outputs provide a rich information base for subsequent signal analysis and processing.
@@
The resulting frequency-domain signal $\mathbf{F}$ has dimensions with $[B,C,F]$, where $F$ is the dimension of the frequency domain. For each band $b\in\{\delta,\theta,\alpha,\beta,\gamma\}$, the band mask is computed and band features are extracted as:
@@
where $\text{MLP}_{\text{spectral}}$ is a network comprising two fully connected layers. To satisfy the Lipschitz continuity constraints, the spectral attention scores are normalized:
@@
where $\mathbf{W}_{s,b}$ is the expanded form of the spectral weights.
The signal is transformed back to the time domain using the Inverse Fast Fourier Transform (IFFT) by applying the IFFT to the combined frequency-domain data $\mathbf{X}_{\text{combined}}$ along the last dimension for a time length of T, and then normalization is applied using LayerNorm.
To enhance model stability, spectral normalization is applied to the weights of the linear layers. This involves scaling the weights $\mathbf{W}$ by their spectral norm $\|\mathbf{W}\|_{2}$ and then multiplying by a Lipschitz constant L:
@@ 3.4 Lipschitz Gradient-Constrained Attention Component (LGCA)
Lipschitz continuity regularization enhances neural networks’ robustness and generalization ability while stabilizing the training process [43, 44]. Inspired by these studies, we designed an LGCA module to enhance the model’s stability, robustness, and generalization ability by restricting its Lipschitz constant [45]. The implementation of LGCA is shown in Figure 1c.
@@
where $\mathbf{S}$ denotes the scaled dot-product scores and $d_{h}$ is the dimension of each attention head.
The attention scores $\mathbf{S}$ are normalized to a probability distribution using the Softmax function. This normalization ensures that the scores sum up to 1, making them suitable for use as weights. The result of this normalization is referred to as the attention weights $\mathbf{P}$.
Then, the dropout operation is applied to the attention weights $\mathbf{P}$ to prevent overfitting. This involves randomly setting a subset of the weights to zero during training, which helps in making the model more robust and generalizable. The resulting attention weights after dropout are denoted as $\mathbf{P^{\prime}}$.
Next, the value tensor $\mathbf{V}$ is combined with the attention weights $\mathbf{P^{\prime}}$ through weighted summation, and the multi-head results are merged into the final output through the output projection layer. Let $\mathbf{Y}$ denote the final output:
@@ 3.5 Lipschitz Gradient-Constrained Normalization Component (LGCN)
To enhance the stability and generalization capability of models when processing complex EEG signals, we propose the LGCN, which is detailed as follows:
@@
This step ensures that the gradient changes of $\mathbf{X}_{\text{lip}}$ are bounded by the Lipschitz constant $L$, which improves training stability and enhances the model’s generalization capability when processing complex EEG signals. The implementation of LGCBE is shown in Figure 1d.
3.6 Loss Function
To effectively evaluate the model’s prediction accuracy, cross - entropy is employed as the loss function for analysis. Cross - entropy loss is a crucial tool for quantifying the discrepancy between the model - predicted probabilities and the one - hot encoded labels in multi - class classification tasks. The model - predicted probabilities are represented as a vector $\bm{\hat{y}}=[\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{n}]$, and the one - hot encoded true labels are represented as a vector $\mathbf{y}=[y_{1},y_{2},\dots,y_{n}]$ (with only one element being 1 and the rest 0). The cross - entropy loss function is defined as follows:
$H(\mathbf{y},\bm{\hat{y}})=-\sum_{i=1}^{n}y_{i}\ln\hat{y}_{i}$ (20)
@@ 4.1 Datasets
To comprehensively evaluate the proposed LEL, we utilized three representative EEG-based emotion recognition datasets.
The EAV dataset [46] provides a multimodal resource for affective computing research, combining synchronized EEG, audio, and video recordings from 42 participants engaged in naturalistic conversational interactions. This dataset captures five distinct emotional states—neutral, anger, happiness, sadness, and calmness—and comprises 8,400 annotated records, establishing itself as a pioneering benchmark in emotion recognition studies. Since the EAV dataset adopts two data collection methods, active and passive, in the experiments, except for experiment Section 4.6, which is used to test the effect of LEL on passive data, the active data of the EAV dataset is used in the other experiments.
The FACED dataset [47] advances EER research by leveraging 32-channel electrophysiological recordings from 123 participants exposed to 28 emotionally evocative video stimuli. This comprehensive resource spans nine discrete emotion categories and is further enriched with subjective self-assessment labels for emotional states, alongside multidimensional affective parameters such as arousal and valence.
@@
4.2 Experimental Setup
4.2.1 Data Preprocessing
We conducted the following preprocessing steps on the raw EEG data: The raw EEG data was resampled to 100Hz to reduce data volume while retaining key features. The resampled data was divided into segments of 500 samples each to extract local features and reduce complexity for subsequent analysis. Finally, the data was partitioned into training and testing sets using stratified sampling to ensure consistent sample distribution, thereby enhancing model training stability and generalization.
4.2.2 Model Training
For model training, the following strategies were adopted: The Adam optimizer was used for model training due to its strong convergence and adaptability, which accelerates the training process and improves model learning efficiency and performance. Additionally, ensemble learning was implemented by freezing classifiers within the model and training each individually before combining them using a learnable weight mechanism. This approach improves the model’s recognition accuracy, stability, and adaptability.
4.2.3 Experimental Validation
To evaluate the validity of the research, the following steps were performed: LEL was evaluated on three datasets: EAV, FACED, and SEED. On the EAV dataset, LEL was compared with results from [46] and the EEG performance of AMERL [15]. On the FACED and SEED datasets, LEL was compared with the original benchmark results. Accuracy (ACC) and F1-score (F1) were used as evaluation metrics. Classification results were visualized using confusion matrices, and model performance was analyzed using Receiver Operating Characteristic (ROC) curves. Additionally, t-SNE visualization was employed to analyze the effects of different training iterations and Lipschitz constants on the model.
4.2.4 Validation Process
The validation process in this study includes the following aspects: First, the effects of Lipschitz continuity constraints on model performance were analyzed from three perspectives: training process, ROC curves, and t-SNE visualization. Second, ablation studies were conducted to evaluate the performance of each individual classifier in the LEL framework and compare it with the full model’s performance to assess the effectiveness of the ensemble learning strategy. Third, LEL’s robustness was validated on passive signals from the EAV dataset. Fourth, LEL’s real-time emotion recognition performance was compared with conventional methods. Finally, the relationships between channels corresponding to different emotions were analyzed using chord diagrams to identify unique neural patterns associated with each emotion.